OpenStudy (niahloves):

Solve using the guess-and-test strategy. The math club has 57 members who are in sixth or seventh grades. There are twice as many sixth graders as seventh graders. How many sixth graders are in the club? A. 38 B. 37 C. 19 D. 18

Vocaloid (vocaloid):

ok, I'm not keen on the "guess and test" strategy, but that means, we pick a number of 6th graders and 7th graders to see if it fits the problem so, as an example, let's pick 30 6th graders, and since there are twice as many sixth graders as 7th graders, there must be 15 7th graders 30 + 15 is less than 57, so we need to pick a higher number

Vocaloid (vocaloid):

keep going until you find the right number

OpenStudy (anonymous):

Okay so you have 57 students in all of this club. If you want to find the sixth graders I suggest you take 57/3. Since you will have all 19's add two 19's together to form 38. That is twice as much sixth graders than seventh graders. Now add 19 to 38 you will come of the answer of 57, that was to check your answer. 6th=38 7th=19 all part of the same club.

OpenStudy (anonymous):

D.

OpenStudy (niahloves):

so it's A?

OpenStudy (niahloves):

@linabear4

OpenStudy (texaschic101):

its not really the guess and test strategy...but this is how to find your answer using substitution... There are 57 sixth and seventh graders... a + b = 57 --- a = 6th and b = 7th Twice as many 6th as 7th.. a = 2b now sub 2b in for a in the first equation a + b = 57 2b + b = 57 3b = 57 b = 57/3 b = 19 a + b = 57 a + 19 = 57 a = 57 - 19 a = 38 so there were 38 sixth graders and 19 seventh graders